# Irreducible symplectic varieties with a large second Betti number

### v3

@@ 4.3. The Jacobian fibration
It remains to prove $b_{2}(\overline{\mathcal{J}})\geq 24$. For a cubic fourfold $[X]\in U$, we denote by $\overline{\mathcal{J}}_{X}$ a $\mathbb{Q}$-factorial terminal irreducible symplectic compactification of the fibration $\mathcal{J}^{\circ}(\mathcal{Y}_{V_{n}})\to V_{n}$ associated with $X$ as above. Let $\Lambda$ be a lattice with a quadratic form $q$, such that there exists an isomorphism $\mu\colon(\mathrm{H}^{2}(\overline{\mathcal{J}}_{X},\mathbb{Z})_{\rm tf},q_{\overline{\mathcal{J}}_{X}})\xrightarrow{\cong}(\Lambda,q)$ where $\mathrm{H}^{2}(\overline{\mathcal{J}}_{X},\mathbb{Z})_{\rm tf}:=\mathrm{H}^{2}(\overline{\mathcal{J}}_{X},\mathbb{Z})/\textrm{torsion}$ and $q_{\overline{\mathcal{J}}_{X}}$ is the Beauville–Bogomolov–Fujiki (BBF) form of $\overline{\mathcal{J}}_{X}$ (see e.g. [9, Section 5.1]). Such an isomorphism $\mu$ is called a $\Lambda$-marking of $\overline{\mathcal{J}}_{X}$. Since the family $\overline{\mathcal{J}}_{U}\to U$ is locally trivial, we know that $(\Lambda,q)$ is independent of the choice of $[X]\in U$ by [9, Lemma 5.7]. Let $\widetilde{U}\to U$ be the universal cover in the analytic topology, where $\widetilde{U}$ is a connected complex manifold. Denote by $\tilde{\pi}\colon\overline{\mathcal{J}}_{\widetilde{U}}\to\widetilde{U}$ the base change of $\overline{\mathcal{J}}_{U}\to U$ to the universal cover $\widetilde{U}$. Then the local system $R^{2}\tilde{\pi}_{*}\mathbb{Z}$ is trivial on $\widetilde{U}$, which implies that $\tilde{\pi}$ is a locally trivial flat family of $\mathbb{Q}$-factorial terminal irreducible symplectic varieties with a $\Lambda$-marking. Let
$\Omega_{\Lambda}:=\{[\sigma]\in\mathbb{P}(\Lambda_{\mathbb{C}})\mid q(\sigma)=0,~{}q(\sigma,\bar{\sigma})>0\}$
be the corresponding period domain of $(\Lambda,q)$ (cf. [9, Definition 8.1]). Then by [9], the family $\tilde{\pi}\colon\overline{\mathcal{J}}_{\widetilde{U}}\to\widetilde{U}$ defines a holomorphic period map
$J\colon\widetilde{U}\to\Omega_{\Lambda}$
by mapping a cubic fourfold $[X]\in U$ with a $\Lambda$-marking $\mu$ on $\overline{\mathcal{J}}_{X}$ to the corresponding period point of $(\overline{\mathcal{J}}_{X},\mu)$. Indeed, $J=\mathfrak{p}\circ J^{\prime}$ is the composition of a period map $\mathfrak{p}\colon\overline{\mathfrak{M}}_{\Lambda}\to\Omega_{\Lambda}$ and a moduli map $J^{\prime}\colon\widetilde{U}\to\overline{\mathfrak{M}}_{\Lambda}$, where $\overline{\mathfrak{M}}_{\Lambda}$ denotes the Hausdorff quotient of the analytic coarse moduli space of $\Lambda$-marked primitive symplectic varieties (see [9, Section 6.12]). Note that $J$ is well-defined on $\widetilde{U}$, as any two $\mathbb{Q}$-factorial terminal irreducible symplectic compactifications of $\mathcal{J}^{\circ}(\mathcal{Y}_{V_{n}})\to V_{n}$ are birational, hence have the same period point by [9, Corollary 6.17].
Let $\mathbf{A}_{21}$ be the moduli space of principally polarized abelian varieties of dimension $21$ and $\mathbf{M}_{5}$ the moduli space of smooth cubic fivefolds. Then there is a period map
@@
given by $\mathscr{P}([Y])=\mathcal{J}(Y)$, which is generically finite onto its image by the local Torelli theorem.
Following the argument in [43, Section 7.4], it is enough to prove $\dim(\operatorname{im}(J))=20$, as this together with the fact that $\rho(\overline{\mathcal{J}}_{X})\geq 2$ implies $\dim(\Omega_{\Lambda})\geq 22$, i.e. $b_{2}(\overline{\mathcal{J}}_{X})\geq 24$. By the local Torelli theorem [9, Proposition 5.5], this reduces to showing that $\dim(\operatorname{im}(J^{\prime}))=20$, i.e. $J^{\prime}$ has generic discrete fibers. If there is an analytic curve $\widetilde{C}\subset\widetilde{U}$ contracted by $J^{\prime}$ passing through a general point of $\widetilde{U}$, then for each pair of general points $[X],[X^{\prime}]\in C$ where $C\subset U$ is the image of $\widetilde{C}$, we have a birational equivalence $\overline{\mathcal{J}}_{X}\dasharrow\overline{\mathcal{J}}_{X^{\prime}}$ by [9, Theorem 6.14]. Since there are at most countably many isotropic line bundles in $\operatorname{Pic}(\overline{\mathcal{J}}_{X})\cong\operatorname{Pic}(\overline{\mathcal{J}}_{X^{\prime}})$, there is an infinite set of points in $C$ such that for $[X]$ and $[X^{\prime}]$ in this set, the following diagram is commutative
@@
is generically finite by [26, Chapter 1, Proposition 5.18], which is a contradiction. ∎
@@
We expect $b_{2}(\overline{\mathcal{J}})=24$, for which one needs to show the image of $J$ has codimension $2$ in $\Omega_{\Lambda}$, in particular, $\rho(\overline{\mathcal{J}})=2$ for a very general $X$. This requires that there exists a smooth or at least terminal relatively minimal compactification of $\mathcal{J}^{\circ}(\mathcal{Y}_{V_{n}})$ over $V_{n}$ with irreducible fibers over codimension one points. In the case of cubic threefolds, using degenerations of Prym varieties, it is shown that the relative compactified Jacobian $\mathcal{J}(\mathcal{Y}_{V_{n}})$ gives such a smooth compactification over $V_{n}$ (cf. [40, Lemma 5.2]). In this paper, we will not address these issues.

### v5

@@ 4.3. The Jacobian fibration
It remains to prove $b_{2}(\overline{\mathcal{J}})\geq 24$. We present two different arguments here. Recall that $\mathrm{H}^{2}(\overline{\mathcal{J}},\mathbb{Q})$ is a pure Hodge structure of weight two with $\mathrm{H}^{2,0}(\overline{\mathcal{J}})=\mathrm{H}^{0}(\overline{\mathcal{J}},\Omega^{[2]}_{\overline{\mathcal{J}}})$ (cf. [51, Theorem 8]). Then the correspondence
$\mathcal{C}\in{\rm CH}^{3}(\mathcal{J}^{\circ}(\mathcal{Y}_{V_{n}})\times X)_{\mathbb{Q}}$
extended from $Z\circ\mathcal{Z}\in{\rm CH}^{3}(\mathcal{J}^{\circ}(\mathcal{Y}_{V_{s}})\times X)_{\mathbb{Q}}$ induces a morphism of rational Hodge structures
$\mathcal{C}^{*}\colon\mathrm{H}^{4}(X,\mathbb{Q})\to\mathrm{H}^{2}(\mathcal{J}^{\circ}(\mathcal{Y}_{V_{n}}),\mathbb{Q})=\mathrm{H}^{2}(\overline{\mathcal{J}},\mathbb{Q})$
that maps $\mathrm{H}^{3,1}(X)=\mathbb{C}\eta$ into $\mathrm{H}^{2,0}(\overline{\mathcal{J}})=\mathbb{C}\overline{\sigma}$. Therefore, the transcendental lattice $\mathrm{H}^{4}(X,\mathbb{Q})_{\mathrm{tr}}$ has nonzero image under $\mathcal{C}^{*}$, and the simplicity of the Hodge structure $\mathrm{H}^{4}(X,\mathbb{Q})_{\mathrm{tr}}$ implies that $\mathcal{C}^{*}$ induces an embedding $\mathrm{H}^{4}(X,\mathbb{Q})_{\mathrm{tr}}\hookrightarrow\mathrm{H}^{2}(\overline{\mathcal{J}},\mathbb{Q})_{\mathrm{tr}}$. When $X$ is very general, we know that $\dim_{\mathbb{Q}}\mathrm{H}^{4}(X,\mathbb{Q})_{\mathrm{tr}}=22$, which gives $\dim_{\mathbb{Q}}\mathrm{H}^{2}(\overline{\mathcal{J}},\mathbb{Q})_{\mathrm{tr}}\geq 22$. Since $\overline{\mathcal{J}}$ has Picard number at least $2$, we obtain $b_{2}(\overline{\mathcal{J}})\geq 24$.
Alternatively, we can also prove $b_{2}(\overline{\mathcal{J}})\geq 24$ by using the geometry of moduli spaces. For a generic cubic fourfold $X$, we denote by $\overline{\mathcal{J}}_{X}$ a $\mathbb{Q}$-factorial terminal irreducible symplectic compactification of the fibration $\mathcal{J}^{\circ}(\mathcal{Y}_{V_{n}})\to V_{n}$ associated with $X$ as above. We fix a $\Lambda$-marking of $\overline{\mathcal{J}}_{X}$, where $\Lambda$ is a lattice, and let $\Omega_{\Lambda}$ be the corresponding period domain (cf. [9, Definition 8.1]). We denote by $\mathbf{M}_{4}$ the $20$-dimensional moduli space of smooth cubic fourfolds. Then the existence of $\overline{\mathcal{J}}_{X}$ defines a rational map
$J\colon\mathbf{M}_{4}\dashrightarrow\Omega_{\Lambda}$
by mapping a general cubic fourfold $[X]\in\mathbf{M}_{4}$ to the corresponding period point of $\overline{\mathcal{J}}_{X}$. Note that $J$ is well-defined, as any two $\mathbb{Q}$-factorial terminal irreducible symplectic compactifications of $\mathcal{J}^{\circ}(\mathcal{Y}_{V_{n}})\to V_{n}$ are birational, hence have the same period point by [9, Corollary 6.17].
Let $\mathbf{A}_{21}$ be the moduli space of principally polarized abelian varieties of dimension $21$ and $\mathbf{M}_{5}$ be the moduli space of smooth cubic fivefolds. Then there is a period map
@@
given by $\mathscr{P}([Y])=\mathcal{J}(Y)$, which is generically finite onto its image by the local Torelli theorem (cf. [16] or [54, Section 6.3.2]).
Following the argument in [43, Section 7.4], it is enough to prove $\dim(\operatorname{im}(J))=20$, i.e. $J$ is generically finite, as that implies $\dim(\Omega_{\Lambda})\geq 21$, i.e. $b_{2}(\overline{\mathcal{J}})\geq 24$. If there is a curve $C\subset\mathbf{M}_{4}$ passing through a general point contracted by $J$, then for each pair of general points $[X],[X^{\prime}]\in C$, we have a birational equivalence $\overline{\mathcal{J}}_{X}\dasharrow\overline{\mathcal{J}}_{X^{\prime}}$ by [9, Theorem 6.14, Theorem 1.1(4)]. Since there are at most countably many isotropic line bundles in $\operatorname{Pic}(\overline{\mathcal{J}}_{X})\cong\operatorname{Pic}(\overline{\mathcal{J}}_{X^{\prime}})$, there is an infinite set of points in $C$ such that for $[X]$ and $[X^{\prime}]$ in this set, the following diagram is commutative
@@
is generically finite by Proposition 4.18, which is a contradiction. ∎
Proposition 4.18.
The rational map $M(\frac{1}{2}-\epsilon)\dashrightarrow\mathbf{M}_{5}\times\mathbf{M}_{4}$ defined by $[(Y,H)]\mapsto([Y],[Y\cap H])$ is generically finite.
Proof.
Let $M^{\mathrm{sm}}$ be the dense open subscheme of $M(\frac{1}{2}-\epsilon)$ parameterizing pairs $(Y,H)$ such that both $Y$ and $Y\cap H$ are smooth. Note that $M^{\mathrm{sm}}$ is open as smoothness is an open condition and that each such pair $(Y,H)$ is $(\frac{1}{2}-\epsilon)$-GIT stable by Proposition 4.7. Then the rational map restricts to a morphism $\Phi\colon M^{\mathrm{sm}}\to\mathbf{M}_{5}\times\mathbf{M}_{4}$. It suffices to show that for a general point $[(Y,H)]\in M^{\mathrm{sm}}$, the set $\Phi^{-1}(\Phi([(Y,H)]))$ is finite. From the definition of the GIT quotient, we may assume that both $Y$ and $H$ are general, and $Y$ has a trivial automorphism group. It is clear that $\Phi_{1}=\mathrm{pr}_{1}\circ\Phi\colon M^{\mathrm{sm}}\to\mathbf{M}_{5}$ is the forgetful map $[(Y,H)]\mapsto[Y]$, where $\mathrm{pr}_{1}\colon\mathbf{M}_{5}\times\mathbf{M}_{4}\to\mathbf{M}_{5}$ is the projection. Then by [26, Chapter 1, Proposition 5.18], we know that the rational map $(\mathbb{P}^{6})^{*}\dashrightarrow\mathbf{M}_{4}$ defined by $H\mapsto[Y\cap H]$ is generically finite. In particular, $f\colon\Phi_{1}^{-1}([Y])\to\mathbf{M}_{4}$ defined by $[(Y,H)]\mapsto[Y\cap H]$ is generically finite as $\Phi_{1}^{-1}([Y])\subset(\mathbb{P}^{6})^{*}$ is a dense open subset by Bertini’s theorem and the triviality of $\operatorname{Aut}(Y)$. Clearly, $\Phi^{-1}(\Phi([(Y,H)]))=f^{-1}([Y\cap H])$. Since $H$ is a general hyperplane in $\mathbb{P}^{6}$, we know that $f^{-1}([Y\cap H])$ is finite, which implies the finiteness of $\Phi^{-1}(\Phi([(Y,H)]))$. The proof is finished. ∎
@@
We expect $b_{2}(\overline{\mathcal{J}})=24$, for which one needs to show the locus of $\overline{\mathcal{J}}$ has codimension $2$ in the corresponding period domain, in particular, $\rho(\overline{\mathcal{J}})=2$ for a very general $X$. This requires that there exists a smooth or at least terminal relatively minimal compactification of $\mathcal{J}^{\circ}(\mathcal{Y}_{V_{n}})$ over $V_{n}$ with irreducible fibers over codimension one points. In the case of cubic threefolds, using degenerations of Prym varieties, it is shown that the relative compactified Jacobian $\mathcal{J}(\mathcal{Y}_{V_{n}})$ gives such a smooth compactification over $V_{n}$ (cf. [40, Lemma 5.2]). In this paper, we will not address these issues.